OpenStudy (anonymous):

find the total of each price item, with working. video game:$49, tax :6.5%

OpenStudy (anonymous):

49 * 1.065

OpenStudy (anonymous):

how did you get the 1.065 ?

OpenStudy (anonymous):

(1 + 6.5%) = 1 + 0.065 = 1.065

OpenStudy (anonymous):

106.5% = 1.056

OpenStudy (anonymous):

ok i get it now